You have 20 stampes that have a total value of $6.62. Some of the stamps are worth $0.24 each and others are worth $0.37 each. How many stamps of each type do you have? Please show how to solve this question.
Found 2 solutions by funmath, checkley71:
Answer by funmath(2933)

You have 20 stampes that have a total value of $6.62. Some of the stamps are worth $0.24 each and others are worth $0.37 each. How many stamps of each type do you have?
It's easier to deal with money in terms of cents to avoid decimals.
Let the number of 24 cent stamps be:x
then because there is 20 stamps all together the number of 37 cent stamps is:20-x
The total value is: 662 cents
:
24x+37(20-x)=662
24x+740-37x=662
-13x+740=662
-13x+740-740=662-740
-13x=-78
-13x/-13=-78/-13
x=6
The number of 24cent stamps: x=6
The number of 37 cent stamps:20-x=20-6=14
:
Check, if you had 6 $.24 stamps and 14 $.37 stamps would the value be $6.62?
6(.24)+14(.37)=6.62 ?
1.44+5.18=6.62 ?
6.62=6.62 !!! We're right!!!
Happy Calculating!!!

Answer by checkley71(8403)

LETX=THE NUMBER OF $.24 CENT STAMPS
LET(20-X)=THE NUMBER OF $.37 CENT STAMPS
.24X+.37(20-X)=6.62
.24X+7.40-.37X=6.62
-.13X=6.62-7.40
-.13X=-.78
X=-.78/-.13
X=6 $.24 CENT STAMPS THEN
20-6=14 $ .37 CENT STAMPS
PROOF
.24*6+.37*14=6.62
1.44+5.18=6.62
6.62=6.62